OpenStudy (anonymous):

9 over 20 subtracted by 2 over 15

OpenStudy (anonymous):

9/20-2/15=19/60

OpenStudy (hoblos):

= 27/60 - 8/60 = 19/60

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

:)